### 17 Dec 2023 v3

@@ 6.1 Simulation Results
All models were generally unbiased or nearly unbiased across all scenarios, and as expected, the biases/RMSEs were smaller and the coverage probabilities (CPs) were closer to their true values for larger sample sizes; see Table 1 and Table A1 for details.
The soft BART models using the full set of predictors, had generally the smallest RMSEs, and CPs close to the nominal values, with an advantage for L-DART. However, under scenario B and C (a strong association between predictors at the last time point while the association for the history is weak or zero), L-DART had much smaller RMSEs compared to S-DART. Moreover, in B and C for the larger sample sizes, as expected, the soft BART model using only the last time-point predictors (S-DART-lag1) performed best in terms of RMSEs and CPs, but by small margin over L-DART. For the smaller sample size (n=100), all CPs $\leq 0.91$ for S-DART-lag1. For all other scenarios, S-DART-lag1 produces RMSEs that are notably larger than the other models and with CPs well below the nominal level, $0.72-0.86$.
Moreover, for all scenarios, the RMSEs for the hard BART models were smaller when fitting 100 regression trees compared to 10, while producing overly conservative credible intervals. In contrst, when fitting 10 trees, CPs were closer to their true values. No improvement in RMSEs and CPs were generally seen for the soft BART models by increasing the number of trees, and the results did not differ significantly from fitting soft BART models with only one tree (Table A2).
@@ 7 Analysis of the Betula data
We consider an intervention where we assign sBP by following the incremental threshold intervention described in Sections 2 and 3. Here, if a subjects’s NVT at times $t=2,\ldots,T$ is greater than $\tau=140$, then sBP is assigned by imposing a incremental shift on the NVT, and otherwise we do not intervene on this subjects. Moreover, we specify a triangular prior for the shift parameter $\delta_{t}$, such that $\delta_{t}\sim Triangular(L_{\tau},0,L_{\tau})$. The lower bound are given by; $L_{\tau}\approx 44$ mm Hg, corresponding to two standard deviations of baseline sBP.
@@
To obtain an estimate of the $SAIE$, we incorporate a prior on the sensitivity parameters in Assumptions C5 and C6. We again specify triangular priors for the sensitivity parameters $\Delta$ and $\lambda$. We let $\Delta\sim\mathrm{Triangular}(0,U_{\Delta},U_{\Delta}),$ with upper bound (and mode) set to $4.6\%$. Here, $4.6\%$ corresponds to the estimated difference in baseline memory between those who are alive at time point $T$ and those who are dead, when adjusting for age and sex; it is unlikely the difference between the strata would to be bigger than this. We further let $\lambda\sim Triangular(0.5,1,1)$, where the upper bound corresponds to a deterministic Monotonicity assumption while the lower bound reflects a moderate correlation between survival under either regimes.
7.1 Main results
@@
We estimated the $PCIE$ and the $SAIE$ using L-DART with priors on the sensitivity parameters as described in the previous section. For each chain we used a thinning of 100, the first 300 iterations were discarded as burn-in, and 2,400 posterior samples were kept. We sampled pseudo data of size $N^{*}=10,000$ at each iteration for the MC integration. Convergence of the posterior samples was monitored using trace plots. Note, all results are given in terms of % correct scores of the maximum (i.e. 76).
@@
Results for the different age-cohorts are given in Figure 1. The results revealed that posterior mean episodic memory score decreased with advancing age both under the natural course and under the hypothetical intervention, although, to a lesser extent for the latter. Among subjects 70 and older, this shift corresponds to postponing memory decline approximately $1.5$ years on average, although the effect is non-significant. The results for the $PCIE$ suggests that there is no long-term cognitive effects of monitoring sBP in this population (credible intervals cover 0 for all age-cohorts), although, the effect is notably stronger for the older age-cohorts compared to middle-aged. The estimates of the $PCIE$ and the $SAIE$ were similar. For the $SAIE$ this suggests that there is no long-term cognitive effects among those surviving irrespective of intervention. Hence, the non-significant results for the $PCIE$ are not driven by altered mortality rates. We moreover note that only small differences were found when comparing survival probabilities under the natural course and under the hypothetical intervention for all age cohorts, and as a result, leading to rather small differences between the $PCIE$ and the $SAIE$.
7.2 Sensitivity analyses
Positivity violations: To assess robustness of results from positivity violations (C3) we discarding subjects with insufficient overlap. To identify these subjects we implemented a generalized propensity score approach, where the propensity score models at times $t=2,\ldots,T$ were fitted using L-DART. Subjects whose sBP was above the given threshold were excluded if their generalized propensity score were out of the range of the generalized propensity score among subjects with sBP within optimal levels.
Missingness for the outcome: Previous studies indicate that the dropout may be MNAR for the outcome [16]. As such, we introduce a sensitivity parameter $\gamma_{t}$ to identify the distribution for dropouts among survivors, to allow for deviations from the MARS assumption (C4). We assume $p(y_{t}|\bar{a}_{t},\bar{l}_{t},r_{t}=0,\bar{s}_{t}=1,\bar{y}_{t-1})=p(y_{t}-\gamma_{t}|\bar{a}_{t},\bar{l}_{t},\bar{r}_{t}=1,\bar{s}_{t}=1,\bar{y}_{t-1})$ for all $t>0$. Our prior belief is that $\gamma_{t}<0$, reflecting a negative shift in memory after a subject drops out. Here, the prior is specified as $\gamma_{t}\sim Triangular(L_{\gamma_{t}},0,L_{\gamma_{t}}),$ where we assume the lower bound not to be bigger than the estimated difference in memory change from baseline to follow-up between dropouts and completers, when adjusting for age and sex. That is, $E[EMS_{i2}-EMS_{i1}|age_{i1},sex_{i1},r_{i3}=1]-E[EMS_{i2}-EMS_{i1}|age_{i1},sex_{i1},r_{i3}=0]$, where $EMS_{it}$ is the episodic memory score at time $t$.
@@ 7.2 Sensitivity analyses
Results: To ease presentation we divided the sample into two age groups; middle-aged, i.e. $age_{iT}<65$, and a older group, i.e $age_{iT}\geq 65$. First, the results were consistent with the results obtained using default assumptions for both age groups when taking into account a unmeasured time-varying confounder, i.e. cholesterol levels. We further investigated the impact of possible positivity violations on the results. For this analysis, subjects with limited overlap ($N=61$) was first discarded. The results did not differ significantly from the main analysis either. Last, the posterior mean episodic memory score was attenuated both under the natural course and under the intervention when using the MNARS assumption instead of the MARS on the outcome. However, the estimates of the $PCIE$ did not change. See detailed results in Table A4.
@@ 8 Conclusions
Using data from a prospective cohort study on aging, cognition and dementia, we implemented a hypothetical incremental threshold intervention to investigate whether blood pressure monitoring at more optimal levels over 15 years was associated with memory improvements. In our study, there were no significant prognostic- or etiological effects across all ages. These findings provide important information about long-term blood pressure control in cognitive aging, indicating that sBP interventions likely do not have a strong effect on memory neither at the population level nor at the individual level. However, more research from long-term clinical trials on whether lowering sBP can prevent cognitive decline is needed to draw firm conclusions.
[…]
We propose a Bayesian semi-parametric approach for the proposed incremental intervention, extending prior research in several ways. First, our approach introduces an incremental intervention for prospective cohort data with attrition, offering a viable alternative when clinical trials are not available or feasible to conduct. second, we offer two types of mortal cohort inferences, providing both prognostic and etiological effect estimates, both of which are critical for public health practice, especially in older populations. Last, our modeling approach employs soft BART models with a novel sparsity-inducing hyperprior for longitudinal data, promoting parsimony in longitudinal prediction settings and demonstrating good performance in simulations.
@@ Acknowledgments
The Supplementary Appendices A and B provide details for the extended G-formula and identification of the SAIE. Computational details for the longitudinal Dirichlet prior and the G-computation algorithm are given in Appendix C. Supplementary simulation results are given in Appendix D and supplementary Betula results are given in Appendix E.
@@ B. Identification of the SAIE under a hypothetical intervention based on the natural value of treatment
where $U=\min\left\{1,\frac{\psi^{g^{int^{*}}}}{\psi^{g^{nc}}}\right\}$. Identification of [S0.Ex14] is obtained by using the law of total probability (ltp) and some algebra. First, for the intervention regime $g^{int^{*}}$, and similarly for $g^{nc}$, we have that
@@
Hence, it follows that (A.4) can be estimated using the g-formula conditioning on survival at a specific time point as described in (S0.Ex2) in the main text.
@@
Hence, it follows that (A.5) can be estimated using the extended g-formula conditioning on survival at a specific time point as described in (S0.Ex7) in the main text. Moreover, by combining the results above we show that the SAIE become (A.2).
@@ C. Computational details
[…]
@@
    where $U=\min\left\{1,\frac{\hat{\psi}^{g^{int^{*}}}}{\hat{\psi}^{g^{nc}}}\right\}$, and one must also sample one set of the sensitivity parameters $\Delta$ and $\lambda$.
@@
D. Simulation results
Table A1: Results in terms of % bias from the simulation study under the different scenarios; A-E using either 1/10/100 regression trees. For scenario A: $\bar{c}=(1,0.75,0.5,0.25)$, B: $\bar{c}=(1,0.75,0.5,0.25)^{3}$, C: $\bar{c}=(1,0,0,0)$, D: $\bar{c}=(1,0.75,0.5,0.25,0,0,0,0)$, and E: $\bar{c}=(0.25,0.25,0.25,0.25)$. We compare five different BART models; L-DART: our approach, H-BART: the standard BART algorithm, H-DART: BART with the sparsity-inducing Dirichlet prior, S-DART: the standard soft BART algorithm, and S-DART-lag1: SoftBart incorporating only the predictors from the last time point.
A B C D E
n=1500
L-DART-1 -0.0 -0.1 -0.0 -0.1 -0.2
S-DART-1 0.0 -0.1 -0.1 -0.1 -0.2
S-DART_l1-1 -0.1 -0.1 -0.0 -0.0 -0.1
H-DART-1 -0.1 -0.1 -0.0 0.0 -0.1
H-BART-1 0.0 -0.0 -0.1 0.0 -0.1
L-DART-10 -0.1 -0.0 0.2 -0.0 -0.1
S-DART-10 -0.1 0.0 0.2 -0.0 -0.1
S-DART_l1-10 -0.0 -0.1 0.2 0.0 -0.1
H-DART-10 -0.1 0.0 0.2 -0.0 -0.0
H-BART-10 -0.1 -0.0 0.2 0.0 -0.1
L-DART-100 -0.1 -0.2 -0.0 -0.1 0.2
S-DART-100 -0.1 -0.2 -0.0 -0.0 0.2
S-DART_l1-100 -0.0 -0.2 -0.0 -0.0 0.1
H-DART-100 -0.0 -0.1 0.0 -0.0 0.2
H-BART-100 -0.0 -0.1 0.1 -0.0 0.2
n=500
L-DART-1 -0.2 -0.2 0.1 -0.2 0.0
S-DART-1 -0.2 -0.1 0.1 -0.2 0.0
S-DART_l1-1 0.0 -0.2 0.1 -0.2 0.0
H-DART-1 -0.0 -0.0 0.3 0.1 0.1
H-BART-1 0.0 0.1 0.2 0.0 0.2
L-DART-10 -0.1 -0.1 -0.0 0.1 -0.1
S-DART-10 -0.1 -0.2 0.0 0.0 -0.1
S-DART_l1-10 -0.1 -0.0 0.1 0.1 0.0
H-DART-10 0.1 -0.0 0.2 0.2 -0.1
H-BART-10 0.1 0.1 0.3 0.1 -0.0
L-DART-100 -0.0 -0.1 -0.3 -0.1 0.1
S-DART-100 -0.0 0.0 -0.3 -0.1 0.1
S-DART_l1-100 0.0 -0.1 -0.2 -0.1 0.1
H-DART-100 0.0 0.1 -0.1 -0.0 0.1
H-BART-100 0.0 0.1 -0.1 -0.0 0.2
n=100
L-DART-1 -0.3 0.1 -0.9 -0.8 -0.5
S-DART-1 -0.3 0.0 -0.6 -0.7 -0.5
S-DART_l1-1 0.0 0.1 -0.7 -0.4 -0.6
H-DART-1 0.2 0.6 -0.0 -0.1 -0.4
H-BART-1 0.5 0.6 -0.0 -0.3 -0.4
L-DART-10 -0.5 -0.2 -0.9 -0.5 -1.1
S-DART-10 -0.5 -0.1 -0.5 -0.5 -1.3
S-DART_l1-10 -0.4 -0.1 -0.8 -0.2 -0.8
H-DART-10 0.2 0.4 -0.5 0.1 -1.0
H-BART-10 0.2 0.5 -0.8 -0.0 -1.0
L-DART-100 -0.6 -0.7 -0.2 -0.4 -0.6
S-DART-100 -0.6 -0.6 0.1 -0.4 -0.6
S-DART_l1-100 -0.1 -0.5 -0.3 -0.3 -0.3
H-DART-100 0.1 -0.2 0.2 0.0 -0.2
H-BART-100 0.0 -0.3 0.2 0.1 -0.1
Table A2: Results from the simulation study under the different scenarios; A - E, of BART models fitted using 1 regression tree. For scenario A: $\bar{c}=(1,0.75,0.5,0.25)$, B: $\bar{c}=(1,0.75,0.5,0.25)^{3}$, C: $\bar{c}=(1,0,0,0)$, D: $\bar{c}=(1,0.75,0.5,0.25,0,0,0,0)$, E: $\bar{c}=(0.25,0.25,0.25,0.25)$. The results are presented in terms of RMSE and 95% coverage probabilities (CP). We compare five different BART models; L-DART: our approach, H-BART: the standard BART algorithm, H-DART: BART with the sparsity-inducing Dirichlet prior, S-DART: the standard soft BART algorithm, and S-DART-lag1: SoftBart incorporating only the predictors from the last time point.
A B C D E
RMSE CP RMSE CP RMSE CP RMSE CP RMSE CP
n=1500
L-DART-1 2.67 0.95 2.05 0.95 1.52 0.95 2.77 0.95 1.68 0.95
S-DART-1 2.69 0.95 2.09 0.95 1.60 0.95 2.79 0.95 1.69 0.95
S-DART_l1-1 5.43 0.86 2.69 0.93 1.50 0.95 5.43 0.86 2.69 0.75
H-DART-1 6.90 0.78 4.78 0.80 3.64 0.81 6.92 0.77 2.89 0.70
H-BART-1 6.98 0.77 4.88 0.78 3.73 0.80 7.01 0.74 2.89 0.72
n=500
L-DART-1 3.59 0.95 2.78 0.95 2.13 0.95 3.80 0.94 2.11 0.95
S-DART-1 3.61 0.95 2.86 0.95 2.45 0.95 3.86 0.95 2.15 0.95
S-DART_l1-1 5.67 0.86 3.11 0.92 2.12 0.95 5.68 0.86 2.88 0.71
H-DART-1 7.29 0.75 5.15 0.76 4.08 0.75 7.33 0.72 2.98 0.73
H-BART-1 7.31 0.75 5.16 0.76 4.11 0.74 7.38 0.71 2.98 0.74
n=100
L-DART-1 5.37 0.95 4.25 0.94 3.65 0.94 5.82 0.95 2.93 0.91
S-DART-1 5.42 0.95 4.63 0.93 4.23 0.90 6.07 0.95 2.94 0.91
S-DART_l1-1 6.58 0.85 4.36 0.90 3.57 0.92 6.58 0.85 3.18 0.76
H-DART-1 7.74 0.75 5.69 0.78 4.65 0.81 7.87 0.72 3.13 0.91
H-BART-1 7.72 0.76 5.68 0.79 4.61 0.81 7.88 0.72 3.15 0.91
E. Betula results

### 20 Mar 2024 v4

@@ 6.1 Simulation Results
The simulation results are presented in detail in Table 1. For all scenarios and models, the absolute relative biases and RMSEs are smaller for the larger sample sizes, as expected. Moreover, the coverage probabilities are close to the nominal values for most scenarios, i.e. CPs: $0.91-0.98$. The soft BART models (including LDART) using the full set of predictors, generally have the smallest biases and RMSEs. LDART improves on SDART in all scenarios except one (Scenario A and for n=1500), where SDART is slightly better than LDART. In Scenario B and C, (that have a strong association between predictors at the last time point and the outcome and weak or zero with previous time points), the soft BART model using only the last time-point predictors (SDART-lag1) performs similarly as the other soft BART models in terms of bias, RMSE, and CPs. For the other two scenarios, SDART-lag1 has higher biases and RMSEs, and smaller CPs ($0.91-0.95$). The two hard BART models produces biases and RMSEs that are notably larger than the other models. In addition, the standard hard BART model produces overly conservative credible intervals in many scenarios, i.e. CPs $\geq 0.97$.
@@ 7 Analysis of the Betula data
We consider an intervention where we assign sBP using the incremental threshold intervention described in Sections 2 and 3. If a subjects’s NVT at times $t=2,\ldots,T$ is greater than $\tau=140$, then sBP is assigned by imposing a incremental shift on the NVT, and otherwise we do not intervene. We specify a triangular prior for the shift parameter $\delta_{t}$, such that $\delta_{t}\sim Triangular(L_{\tau},0,L_{\tau})$. The lower bound and mode are given by $L_{\tau}\approx 44$ mm Hg, corresponding to two standard deviations of baseline sBP, and the upper bound is 0. The intervention implies a negative shift from the NVT with between 0 and 44 mm Hg for individuals with a $NVT>140$ at each follow-up time point.
@@
To estimate the $SAIE$, we also specify priors on the sensitivity parameters in Assumptions C5 and C6. We specify triangular priors for the sensitivity parameters $\Delta$ and $\lambda$. For $\Delta$, we assume $\Delta\sim\mathrm{Triangular}(0,U_{\Delta},U_{\Delta}),$ with upper bound (and mode) set to $4.6\%$; $4.6\%$ corresponds to the estimated difference in baseline memory between those who are alive at time point $T$ and those who are dead, adjusting for age and sex. It is unlikely the difference between the strata would be bigger than this. We further assume $\lambda\sim Triangular(0.5,1,1)$, where the upper bound corresponds to the deterministic monotonicity assumption while the lower bound reflects a moderate correlation between survival under the two regimes.
Results
@@
We estimated the $PCIE$ and the $SAIE$ using L-DART with priors on the sensitivity parameters as described above. For each chain we used a thinning of 100, discarded the first 300 iterations as burn-in, and kept 2,400 posterior samples for inference. We sampled pseudo data of size $N^{*}=10,000$ at each iteration for the MC integration. Convergence of the posterior samples was monitored using trace plots. Note, the memory outcome, i.e. a test score, is presented in %.
@@
Results for the different age-cohorts are shown in Figure 1. The results revealed that posterior mean episodic memory score at end of follow-up was attenuated with advancing age, both under the natural course and under the hypothetical intervention, (although to a lesser extent for the latter). Among subjects 70 and older, the shift in memory from the intervention corresponds to postponing memory decline approximately $1.5$ years on average, although the effect is non-significant. The results for the $PCIE$ suggest that there is no long-term cognitive effects of monitoring sBP in this population (credible intervals cover 0 for all age-cohorts), although the effect is notably stronger for the older age-cohorts compared to middle-aged. The estimates of the $PCIE$ and the $SAIE$ were similar. For the $SAIE$ this suggests that there is no long-term cognitive effects in the always survivor stratum, i.e. the stratum where individuals were to survive under both the natural course and the intervention. Hence, the non-significant results for the $PCIE$ are not driven by differential mortality rates. We moreover note that only slightly higher survival probabilities were obtained under the the hypothetical intervention for all age cohorts when compared to survival probabilities under the natural course, and as a result, leading to rather small differences between the $PCIE$ and the $SAIE$.
We performed sensitivity analyses to investigate robustness of results to violations of the assumptions C2-C6. Details are given in Appendix D and results are summarized in Table A2. To ease presentation we divided the sample into two age groups; middle-aged, i.e. $age_{iT}<65$, and a older group, i.e $age_{iT}\geq 65$.
@@
First, the results are not sensitive to other (plausible) specifications of the priors on $\lambda$ and $\Delta$, data not shown. Moreover, the results were consistent with the results obtained using default assumptions for both age groups when taking into account a unmeasured time-varying confounder. We further investigated the impact of possible positivity violations on the results. For this analysis, 61 subjects with limited overlap were discarded. The results did not differ significantly from the main analysis, although the effect was somewhat attenuated for both middle aged and older. Using the missing not at random (MNARS) assumption instead of the MARS for the outcome resulted in attenuated posterior mean episodic memory score for both regimes. However, the estimates of the $PCIE$ were stable. Finally, the results did not differ significantly from the main analysis when fitting the two other soft BART models, i.e. S-DART and S-DART-lag1, instead of LDART. However, the effects were much smaller for both middle aged and older.
@@ 8 Conclusions
Using data from a prospective cohort study on aging, cognition and dementia, we investigated population effects on episodic memory of an hypothetical incremental threshold intervention over 15 years that reduced sBP in individuals with hypertension. In our study, there were no significant prognostic- or etiological effects, although effects were markedly stronger in older individuals. These findings provide important information about long-term blood pressure control in cognitive aging, indicating that sBP interventions likely do not have a strong effect on memory, whether at the overall population level or in the population that would survive under both the natural course and the intervention. However, more research from long-term clinical trials on whether lowering sBP can prevent cognitive decline is needed to draw firm conclusions.
We propose a Bayesian semi-parametric approach for the proposed incremental threshold intervention, extending prior research in several ways. First, our approach introduces an incremental intervention for prospective cohort data with attrition, offering a viable alternative when clinical trials are not available or feasible to conduct. Second, we offer two types of mortal cohort inferences, providing both prognostic and etiological effect estimates, both of which are critical for public health practice, especially in older populations. Last, our modeling approach employs soft BART models with a novel sparsity-inducing hyperprior for longitudinal data, promoting parsimony in longitudinal prediction settings and demonstrating good performance in simulations.
Concerns have been raised about principal stratification methods relying on assumptions that are often hard to justify in practice (e.g. [5]). Monotonicity is one such assumption. Our approach relies on the less strong stochastic monotonicity assumption and uses sensitivity parameters that allow researcher to investigate sensitivity of results to this unverifiable assumption. However, a limitation with proposed principal stratification approach is that we only included age as a survival-predictive covariate for partial identification of the SAIE. In future work we will explore the possibility of incorporating more baseline covariates (see for example [29, 34, 35]).
@@ Acknowledgments
The Supplementary Appendices A and B provide details for the extended G-formula and identification of the SAIE. Computational details for the longitudinal Dirichlet prior and the G-computation algorithm are given in Appendix C. Supplementary Betula results are given in Appendix D.
@@ B. Identification of the SAIE under a hypothetical intervention based on the natural value of treatment
where $U=\min\left\{1,\frac{\psi^{g^{int^{*}}}}{\psi^{g^{nc}}}\right\}$. Identification of [S0.Ex10] is obtained by using the law of total probability (ltp) and some algebra. First, for the intervention regime $g^{int^{*}}$, and similarly for $g^{nc}$, we have that
@@
Hence, it follows that (A.4) can be estimated using the g-formula conditioning on survival at a specific time point as described in (S0.Ex1) in the main text.
@@
Hence, it follows that (A.5) can be estimated using the extended g-formula conditioning on survival at a specific time point as described in (S0.Ex3) in the main text. Moreover, by combining the results above we show that the SAIE become (A.2).
@@ C. Computational details
Similar to Linero and Yang (2018), we obtain simple conjugate posterior updates for $(w,v^{t},u,v^{k})$ by implementing the Bayesian backfitting approach to sample from the posterior using the Markov chain Monte Carlo algorithm [2]. In particular, the full conditional for $w$ is given by $w\sim Beta\left(a+m_{w_{1}},b+m_{w_{2}}\right)$, where $m_{w_{1}}$ is the number of branch splits for the set of current predictors and $m_{w_{2}}$ is the number of branch splits for all the sets of preceding time-varying covariates. Similarly, the full conditional of $\mathbf{v}^{t}$ and $\mathbf{v}^{k}$ are given by $\mathbf{v}^{t}\sim\mathcal{D}\left(\eta/P_{t}+m_{v^{t}_{1}},\ldots,\eta/P_{t}+m_{v^{t}_{P_{t}}}\right)$ and $\mathbf{v}^{k}\sim\mathcal{D}\left(\eta/P_{k}+m_{v^{k}_{1}},\ldots,\eta/P_{k}+m_{v^{k}_{P_{k}}}\right)$ respectively, where $m_{v^{k}_{j}}$ is the number of branch splits on the $j$th predictor in the set of predictors associated with time point $k=1,\ldots,t-1$. Finally, the full conditional of $\mathbf{u}$ is given by $\mathbf{u}\sim D\left(\frac{\alpha_{1}}{t-1}+m_{u_{1}},\cdots,\frac{\alpha_{t-1}}{t-1}+m_{u_{t-1}}\right)$, where $m_{u_{j}}$ is the number of branch splits on the predictors in the $j$th set of time-varying covariates.
@@
    where $U=\min\left\{1,\frac{\hat{\psi}^{g^{int^{*}}}}{\hat{\psi}^{g^{nc}}}\right\}$, one set of the sensitivity parameters $\Delta$ and $\lambda$ are samples.
@@
D. Supplementary Betula results